16m^2+25. Factor completely


9a^2=3a. solve the equation

Answer by checkley75(3666)

16m^2+25 using the quadratic equation we get:
m=(0+-sqrt[0^2-4*16*25])/2*16
m=(+-sqrt-1600)/32
m=(+-40i)/32
m=40i/32
m=1.25i answer.
m=-40i/32
m=-1.25i answer.
--------------------------------------------------------
9a^2=3a
9a^2-3a=0
a(9a-3)=0
a=0 answer.
9a-3=0
9a=3
a=3/9
a=1/3 answer.